OpenStudy (anonymous):

Im in Intermediate Algebra and I need help. Simplifying radicals using the multiplication property of radicals X^5 under square root . How do I do this problem.

OpenStudy (jlastino):

you can use the law of exponents...since you are asked to find the square root sqrt x^5 would also be equal to x^2 x^2 x (using the law of exponents,when we multiply exponents above the same base, we add them. we know that sqrt x^2 =x sqrt x^5 =x^2 sqrt x

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

thats confusing